OpenStudy (anonymous):

Samantha is going on vacation for the summer and is trying to choose between two different plans. The first plan costs $450 for 3 days at a hotel and 2 days at an amusement park. The second plan offers 5 days at the same hotel and 3 days at the amusement park for $700. The cost of 1 day at the hotel and the amusement park is the same under both plans. How much does a 1 day trip to the amusement park cost? $50 $75 $150 $200

OpenStudy (anonymous):

let the hotel cost = $ x per day and let the cost of amusement park = $ y per day So for first plan 3x + 2y = 450 ------------ (I) and for second plan 5x + 3y = 700 ------------ (II) Multiply (I) by 5 and (II) by -3 and add the resulting equations 15x + 10y = 2250 -15x - 9y = -2100 we get y = 150 So, the cost of 1 day trip to the amusement park is $ 150